OpenStudy (ineedhelpnow):

find the x-intercepts and the y-intercept for the parabola whose equation is given. if an intercept is an irrational number, round it to the nearest tenth. y=-x^2+9x-20

OpenStudy (anonymous):

for the x intercept, set y=0 -(x^2-9x+20)=0 (x-5)(x-4) =0 x=5, x=4 x-intercepts are (5,0) and (4,0) for the y intercept set x=0 y=-20 y intercept is (0,-20)